OpenStudy (anonymous):

Quick Question: x + 0.05x = 250. Solve for x.

OpenStudy (solomonzelman):

\(\large\color{black}{ x + 0.05x = 250 }\) I am going to color the like terms for you: \(\large\color{black}{ \color{red}{x + 0.05x} = 250 }\)

OpenStudy (igreen):

First combine like terms.

OpenStudy (solomonzelman):

can you add the red part?

OpenStudy (solomonzelman):

Hint: \(\large\color{black}{ \color{red}{ x + 0.05x} = 250 }\) is same as: \(\large\color{black}{ \color{red}{ 1.00x + 0.05x} = 250 }\)

OpenStudy (anonymous):

okay got it

OpenStudy (solomonzelman):

got the answer, or how to add the like terms (part in red) ?

OpenStudy (anonymous):

x = 238.1, the decimal just jumbled my brain for a sec.

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

you are correct!